OpenStudy (anonymous):

Enough of a monoprotic acid is dissolved in water to produce a 0.0113 M solution. The pH of the resulting solution is 2.33. Calculate the Ka for the acid.

OpenStudy (australopithecus):

this should help

OpenStudy (cuanchi):

pH= - log [H+] => [H+] = 10^(-pH) \[inonization.percentage= \frac{ [H ^{+}] }{ C _{a} }\times 100\] the % of dissociation of the acid is > 5% we can not assume the x is too small and state the equilibrium use the formula \[Ka=\frac{ [A ^{-} ][H ^{+}]}{ C _{a}-[H ^{+}] }\] [A-]=[H+] \[K _{a}= \frac{ [H ^{+}]^{2} }{ C _{a} -[H ^{+}]}\]